NEWS_2003

These are the news starting 01 Jan 2003
News prior to 2001 have been put to file news_old ; 2001 and 2002 year news are in files news_2001 and news_2002

20 Dec 2003 The new equipment works fine: there remains a problem with the WBGT heat-stress sensor, which is not acute; the solar radiation sensor is not yet installed on the new station, but data are still collected on the old equipment, and will be integrated into the monthly data file later on.
Here are the data channels currently in use: 61 & 62, 1 to 5, 9 to 11, 14 to 22. Please discard all other channels for the moment! Please take note that the minimum and maximum values in the data files should be discarded, as these are global min-max values which may include non-working sensor data!
16Dec 2003 At 16:00 today meteoLCD switched to its new equipments in the new location (bye..bye old faithful container!). Some sensors are not yet working: ESR (ch6, solarrad) is the most notable missing. Some of the older sensors (like the defunct dewpoint, ch12&13) will not be reintroduced. 
Many sensors are brandnew:
Airtemp, Relative Humidity, Windspeed and Winddirection (ultrasonic sensor), ozone (API400)...
All computers have been changed, several are brand-new, other vintage refurbished material .
Please bear some teething problems with me...the 7 days graphs are self-healing !
There (probably) will be no missing data for December, as the move has been carefully planned.
Pictures and more asap!
BIG BIG THANKS to all the people who made this  upgrade possible! 
15 Dec 2003 Barometer removed...it will be installed at the new station today. We plan to move the gas sensors tomorrow, and switch over to the new station in the evening...
11 Dec  2003 Problem repaired ( a NIC died...); live data available again. No data loss.
10 Dec 2003 1. There is a big problem with the main computer...will be repaired tomorrow
2. I moved the UVA sensor (ch.10) to the new location; so UVA live data will be unavailable for about a week (we plan to switch over to the new site at the end of next week...)
07 Dec 2003 Uploaded Nov03 data.
Small change to the total ozone graph (dobson) to correct garbled x-axis
17 Nov 2003 Removed WBGT sensors for migration to new station. Sensors will be unavailable for some time...no heat-stress danger for the moment!
08 Nov 2003 Inserted link to Earthtrack satellite data of globe temperatures (Northern & Southern Hemisphere) in "Today's Data..." page. Removed links to GOME: instrument seems definitively out of order!
01 Nov 2003 Hard disk failure on one of the computers at 30Oct03 00:00; computer replaced today at 09:00. No data lost, but some bugs in the 7days graphs (will vanish...). Added link to near-live picture of solar activity (GOES, NOAA)
19 Oct 2003 Added computation of a more correct total sunshine duration to the BIG Excel file 2002_OKK_fm.xls; this calculation is based on a new procedure developed by Jean Olivieri from Meteo France.
14 Oct 2003 Recalibration of CO2 and NOx sensor
13 Oct 2003 Some editing done of BIG EXCEL file 2002_OKK_fm.zip: only CO2 data in the range [300...600] ppmV have been kept.
10 Oct 2003 Uploaded September 2003 data to the data archive.
03Oct
2003
Maintenance work by Envitec on O3, NOx and CO2 sensors ( see the spikes!)
25 Sep 2003 The IP address of our winnetou.lcd.lu server used by the lightning display has been changed. If you have problems loading the display, please define one of your DNS servers as 158.64.62.138. This is just a stop gap....
11 Sep 2003 1. Lightning computer ok again and running...
2. August data upload (first half of month = heatwave: max.temp = 38.4C on 08Aug 16:00 UTC)
29 Aug 2003 There is a problem with the lightning data (probably a breakdown of a local LAN switch...) Sorry!
23 Aug 2003 A file security problem prevented the correct creation of the plots... sorry for this restarting problem. Solved at 15:23 UTC. 
22 Aug 2003 Live data display available again. One of the harddisks of the main Linux computer had a file problem that was unrepairable during automatic startup. No data have been lost. See the data archive for July data and the "all data since last logger restart" for the data since July 2003 ( available probably from 23Aug03 on).
The live data display starts at today 07:30 UTC.
13 Aug 2003 Live data display is unavailable since 5th Aug. due to a computer failure. This can not be repaired before 22th Aug.
27 Jul 2003 There was an error in the time legend at the bottom of the live graphs (year was 2001 instead of 2003). Corrected.
Logger (nearly full) memory cleared and logger restarted at 14:00:30 UTC
26 Jul 2003 Routine maintenance work (see the spikes!). Lightning page changed to display a larger radar-screen. Added explanations on icons of the different lightning types.
25 Jul 2003 The Boltek Lightning sensor went into limbo on the 21th July after...you guess it: a severe thunderstorm! All is well again starting 16:00 UC
07 Jul 2003 Maintenance work by Envitec on ozone, NOx and CO2 sensors (this explains erratic measurements and spikes)
30 Jun 2003 ...Ooops! The Lightning/2000 software needs a special instruction to enable the hardware; this has been missed during the last restart, so there were no strokes recorded during the previous days. This has been corrected starting today 11:00 UTC.
The html code has been modified to update the browser every minute, freeing the user from hitting the reload button to get a new screen. There was a nice storm passing over Luxembourg this afternoon.
24 Jun 2003 Lightning detector from BOLTEK installed. The system is not yet integrated into meteoLCD, so the link on the home page should be taken as an alpha-release! More on this asap...
31 May 2003 A new paper on the influence of pointing errors on the Microtops II sunphotometer readings.
28 May
2003
Maintenance and recalibration on all gas sensors. The NO2 readings are too high after this. Will be checked asap...
18May
2003
The Pt100 DRYBULB sensor (CH27) seems too low; the computation of the WBGT now uses the AIRTEMP sensor (CH7)
10May
2003
A new round in the fight against the birds picking at the wetbulb sensor: a plastic meshed cover has been constructed and fixed to cover this sensor without restraining (too much) the free air flow. The last day, the wick often lasted only several hours, before a bird tore it away or just moved it out of the water bucket (... transforming the wet sensor into a dry one!). Click here for a picture of the anti-bird-protection!
02 May 2003 1. started total ozone measurements using a new Microtops II sensor. Preliminary intercomparisions show that the old sensor is probably 7% too high, what confirms the comparisons with Uccle and satellite measurements. 
2. April 03 data uploaded to archive.
24 Apr 2003 The WetBulb sensor (channel 21, an element of the WBGT system) has been replaced. This Pt100 sensor has probably been attacked by birds who broke the glass sensor. The Pt100 has been replaced by a Pt1000 sensor in aluminium casing, who hopefully will resist the furious attacks of these flying hackers!
Please take a look at the README file describing the sensors, because the format of data column 21 has changed!
23 Apr 2003 08:00 No ground ozone data today... the sensor is on the move!
16:00 Sensor back and operational again.
11 Apr 2003 Malfunction of WetBulb sensor (probably glasstube with Pt100 broken).This can not be repaired before 2 weeks.
03Apr03 Uploaded March 2003 data
28 Mar 2003 A nasty bug survived in the scripts to make the total ozone plot out of the html file. It has been corrected, together with some tidying up...
24 Mar 2003 Maintenance and recalibration work starting 12:30 UTC. There was a problem with the wetbulb sensor the last 2 days: the wet wick has (probably) been picked out of the water reservoir by thirsty birds, so it dried out. We will keep a sharp eye on these feathered companions (mainly crows) and if necessary fence the sensor off ! :-)
18 Mar 2003 Now plot of total ozone column is DOBSON chapter! God bless the makers of Gnuplot and Perl!
17 Mar 2003 Pt100 sensor of the wetbulb device repaired and recalibrated. This work explains the spikes on the graphs. Logger downloaded and restarted (mandatory as config. file was changed).
15 Mar 2003 The Pt100 sensor of the wet-bulb thermometer is broken; measurements are still ok, but the device will be repaired asap...
09 Mar 2003 A short and extremely simple paper on how to compute the variation of the UVI during the day
(in French, PDF format)
03 Mar 2003 A short HOWTO fetch the global ozone column for Diekirch measured by the Gome instrument (ESR-2 satellite)
25 Feb 2003 Update of the intercomparison on total ozone measurements done at the RMI (Uccle) and meteoLCD
20 Feb 2003 A short report written by students of the LCD comparing two solar sensors (ESR and CM21)
03 Feb 2003 A new short report from Francis Massen on the relationship between the zero-offset of the 501 UVA Biometer and deviation of its ionternal thermal stabilization
02Feb
2003
1. There was small error in the date column of the 2002_OKK_fm.zip big Excel file; this has been corrected.
2. A comparison with wetbulb and air_temperature sensors shows that the first is too low when completely encased in ice ( temperature < -5C). A corrective offset of +1.2C is applied to the graphs, but not yet to  the raw data. We will wait for warmer days to make a comparison in dry conditions before taking more definitive action. Btw, heatstress is a somehow academic problem for the moment!
3. New version of readme.html sensor description file with some updated warnings!
01Feb
2003
Uploaded January 2003 data file. There is a small problem with the thermal stabilization of the UVA_Bio-
meter (which explains too high night readings). Everything back normal from 03Jan03 on...
31 Jan 2003 The counter topped 50000 visitors (since 1996)! Thank you folks!
24 Jan 2003 Recalibration and maintenance work.
19 Jan 2003 Our "roof movie" is updated by 2 new pictures!
17 Jan 2003 I have been told today that our research project meteoSOLAIR is deadborn. Read here about this sad story!
04 Jan 2003 Problems with logging solved: the computer hang during a reboot, and several ( most!) data files were corrupt... A new NEWS file for 2003 will be started now!
Upload of dec02.dat data file of December 2002 data and full 2002 year data zipped file
03 Jan 2003 The system stopped transferring data the 01Jan03 after 23:30 UTC. Hopefully the battery backup is enough to assure continuing logging. Repair will be tried tomorrow morning...

 


Back to Home-Page of Meteo_LCD